# Emergent active turbulence and intermittency in dense algal suspensions of Chlamydomonas reinhardtii

### v2

@@ Abstract
Active-fluid turbulence has been found in bacterial suspensions, but not so far in their algal counterparts. We present the first experimental evidence for turbulence in dense algal suspensions of Chlamydomonas reinhardtii. We carry out a detailed analysis of the statistical properties of the flow present in these cell suspensions and show that they are quantitatively distinct from their counterparts in two-dimensional fluid and bacterial turbulence. Both kinetic-energy and density spectra of the fluid flow in algal turbulence show power-law regimes with unique scaling exponents. The fluid velocity probability distribution function (PDF) is strongly non-Gaussian and the length dependence of the PDF of fluid-velocity increments indicates small-scale intermittency. We compare and contrast our results with recent theoretical predictions for active-scalar turbulence and active glasses. Overall, our results highlight that active turbulence can arise, even in absence of orientational instabilities, so it is not limited to bacterial suspensions but it can also be found in many biological systems with free-swimming micro-organisms.
@@ I Uncovering Algal turbulence
Figure 1: Time-lapse images of the motion of an isolated (a) wild-type (WT) and (b) mbo2 C. reinhardtii cell [scale bar, $10$ microns]. Experimentally measured beat averaged flow-fields of isolated (c) WT and (d) mbo2 C. reinhardtii cells; the arrow indicates the direction of motion. (e) Velocity-vectors, obtained from particle-image velocimetry (PIV), overlaid on the image of a dense suspension of WT C. reinhardtii cells. (f) Optical image of a suspension of WT C. reinhardtii cells, with average density $\bar{\rho}=0.49$ [scale bar, $50$ microns]; pseudocolor plots of (g) the density and (h) the vorticity $\omega$ [colour-bar unit $sec^{-1}$] fields, for the image shown in (f). Panels (i), (j), and (k) are the mbo2 C. reinhardtii counterparts of (f), (g), and (h), respectively. For the spatiotemporal evolution of (f)-(k) see the videos V1-V5 in the Supplementary Information.
@@
Turbulence, which is ubiquitous from astrophysical to cellular scales [1, 2, 3, 4, 5, 6, 7], continues to provide new and exciting challenges for physicists, engineers, and mathematicians. Nonequilibrium turbulence-like states, which have been found over the last decade or so in dense bacterial suspensions [8, 9, 10, 11, 12, 13, 14], provide important recent examples of new types of turbulence. This has been christened active turbulence because such suspensions are active fluids in which energy is injected into the fluid not by an external force, as in conventional fluid turbulence, but by the conversion of chemical sources of energy to kinetic energy by the constituents in the suspension [12].
Our work is inherently interdisciplinary because it studies complex flows in a biological system using techniques from turbulence theories and simulations. We present the first experimental study of emergent active turbulence and intermittency in dense algal systems. In particular, we carry out experiments on suspensions of two types of Chlamydomonas reinhardtii (henceforth, C. reinhardtii), the wild type (WT) and the mutant mbo2. We then characterise the statistical properties of this algal turbulence using measures that are employed to study conventional fluid turbulence. We compare our results with recent theoretical and numerical studies of scalar active turbulence [15].
We obtain kinetic-energy and algal-concentration spectra, longitudinal-velocity structure functions [1, 3, 4, 5, 16], probability distribution functions (PDFs) of velocity components, length-scale dependent longitudinal-velocity increments, and the Okubo-Weiss parameter $\Lambda$, which distinguishes between vortical and extensional regions in a flow. These spectra display power-law scaling regions as a function of the wavenumber $k$; and they indicate that kinetic-energy and concentration fluctuations are spread over a wide range of spatial scales; these are clear signatures of turbulence. However, the power-law exponents that characterise these scaling regions are distinctly different from their fluid-turbulence and bacterial-turbulence counterparts [1, 3, 4, 5, 12], as are the PDFs and structure functions mentioned above. Thus, our investigations uncover a new type of active turbulence, which we call algal turbulence, and whose special statistical properties we elucidate below. Earlier experimental and theoretical studies [17, 18, 19] have investigated enhanced tracer diffusion in dilute suspensions of swimming eukaryotic swimmers like C. reinhardtii. However, this enhanced tracer diffusion is more akin to Lagrangian chaos or passive-scalar intermittency in simple flows [20, 21, 22] than to the fully developed algal turbulence we have uncovered here.
As we increase the concentration of C. reinhardtii cells, the temporal evolution of our dense algal suspensions slows down and the intensity of algal turbulence in our systems decreases. Therefore, it behooves us to explore whether our algal systems cross over from algal turbulence to an active algal glass. To examine this possibility for the cells in our suspensions, we calculate several quantities that are used to characterise slow dynamics in active glasses [23, 24, 25], including (a) the mean-square displacement (MSD), (b) the self-intermediate scattering function $F_{s}(k,t)$, (c) the overlap function $Q(t)$, and (d) the four-point correlation function $\chi_{4}(t)$, which characterises the dynamic heterogeneity in a glassy system. Our systems exhibit dynamic heterogeneity and distinct slowing down as $\bar{\rho}$ increases, but they do not display all the characteristic properties of an active glass.
II Statistical characterisation of dense algal suspensions
@@
Figure 3: Semilog plots of different probability distribution functions (PDFs): PDFs of (a) the $x$-component of the velocity field $\bm{u}$ and (b) the Okubo-Weiss parameter $\Lambda$ [see Eq. (10)] for WT C. reinhardtii cells and different mean densities $\bar{\rho}$. Panels (c) and (d) are the mbo2 C. reinhardtii counterparts of (a) and (b), respectively. Our data are averaged over different samples with similar values of $\bar{\rho}$ (within $\simeq 10\%$ of the mean); the error bars denote ±$\pm\varsigma(\mathcal{P}(u_{x}))$ and ±$\pm\varsigma(\mathcal{P}(\Lambda))$.
Figure 4: Semilog plots, for WT C. reinhardtii cells, of (a) the PDFs of longitudinal-velocity increment of the velocity field $\bm{u}$, for different values of the separation $l$ and the mean density $\bar{\rho}=0.43\pm 0.06$, and (b) the flatness $F_{4}$ [see Eq. (11)] versus $l/a_{w}$ for different values $\bar{\rho}$. The plots in (c) and (d) are the mbo2-mutant counterparts of those in (a) and (b), respectively. Our data are averaged over different samples with similar values of $\bar{\rho}$ (within $\simeq 10\%$ of the mean); the error bars denote ±$\pm\varsigma(\mathcal{P}(\delta\bm{u}_{\parallel}(l)))$ and ±$\pm\varsigma(F_{4})$.
Figure 5: Plots versus the scaled time $\Delta t/\tau_{w}$ of (a) the mean-square displacement (MSD) [the local slope $\frac{\delta(log\langle\Delta x^{2}\rangle)}{\delta(log\Delta t)}$ is shown in the inset], (b) the modulus $|F_{s}(k_{x},t)|$ of the self-intermediate-scattering function (for $k_{x}a_{w}=8.35$), and (c) the self-overlap function $Q(t)$ and the four-point correlation function $\chi_{4}(t)$ [see Eqs.(12), (13), (14), (15)] for WT C. reinhardtii cells with $\bar{\rho}=0.74\pm 0.05$; the reference time scale $\tau_{w}$ is defined in the text; (d), (e) and (f) are the mbo2-mutant counterparts of those in (a), (b) and (c), respectively. Our data are averaged over cell-trajectories of samples with similar values of $\bar{\rho}$ (within $\simeq 10\%$ of the mean); the error bars denote ±$\pm\varsigma(\Delta x^{2}(t))$ and ±$\pm|\varsigma(F_{s}(k_{x},t))|$.
In Fig. 1 we illustrate the two algal systems we consider. Wild type (WT) C. reinhardtii is a contractile swimmer, on beat-averaged time scales and in the far-field limit [Figs. 1(a) and (c)]: it pulls the fluid from its front and back and pushes the fluid out from its sides. The two flagella of WT C. reinhardtii beat in a breast-stroke fashion as they propel the swimmer forward [Fig. 1(a)]. We compare and contrast the active turbulence of WT cells with that of the mutant mbo2 C. reinhardtii, whose cells swim at a significantly lower speed ($v_{m}\simeq$ 50 $\mu$m/s) than those of WT C. reinhardtii, ($v_{w}\simeq$ 100 $\mu$m/s) and are effectively rear propelled by the flagella. The beat-averaged flow field of mbo2 cells resembles that of an extensile swimmer in the far-field limit [Figs. 1(b), (d)]. We focus on collections of such swimmers, confined to a quasi-two-dimensional (2D) domain, wherein the cells swim in a wide chamber with a depth that is comparable to, but slightly larger than, the cell diameter [Materials and Methods]. This allows us to visualize the spatiotemporal evolution of a monolayer of freely swimming C. reinhardtii cells. The swimmers collectively churn the fluid, which leads to an emergent chaotic flow, with more complex vortical fields than those generated by individual swimmers. We characterise this complexity systematically using statistical measures that are employed to analyse statistically homogeneous and isotropic turbulence in fluids (see below). Figure 1(e) shows an optical image of the swimmers, at a representative time, overlaid with velocity vectors of the flow field obtained using PIV, whose magnitude is given by the colour that goes from blue (low speed) to red (high speed). We monitor the distributions of cells in dense suspensions of C. reinhardtii, and show these in Fig. 1 for WT cells (here and henceforth in a blue panel) and the mbo2 mutant (here and henceforth in a beige panel). Figure. 1(f) shows a microscopic snapshot of a WT suspension with an average density $\bar{\rho}=0.49$, at a representative time. At the same instant, we show a pseudocolor plot of the density field $\rho$, which lies between 00 (no cell present) and $1$ (complete coverage with the cell bodies [Fig. 1(g)] and the vorticity-field $\omega$ computed from the velocity field $\bm{u}$ [Fig. 1(h)]. The vorticity plots reveal higher magnitudes of the vorticity for the WT cells in comparioson to their mbo2 variants; this arises principally from the differences in their swimming speeds. Figures 1 (i), (j), and (k) are, respectively, the counterparts of Figs. 1 (f), (g), and (h) for the mbo2 mutant. These plots and the videos V1-V5 show that dense suspensions of both WT and mbo2 C. reinhardtii cells show emergent nonequilibrium states that appear to display spatiotemporal chaos and a type of turbulence. We quantify the statistical properties of these states below.
@@
In Fig. 2 we present log-log plots versus the wavenumber $k$ of the energy and concentration spectra, $E(k)$ and $\Phi(k)$, respectively [Eqs. (6)-(8) in Subsection IV.5], for the wild-type (blue panels) and mbo2 mutant (beige panels). We compensate these spectra with different powers of $k$ to uncover different power-law regimes; and we use $a_{w}$ ($a_{m}$), the mean wild-type (mbo2-mutant) cell size (see Fig. 1 in the Supplementary Information) to non-dimensionalise $k$. The energy spectra scale as $E(k)\sim k^{1/2}$ in the $k\rightarrow 0$ limit for all $\bar{\rho}$ and for both types of swimmers. In the high-wavenumber limit, we observe an interesting transition in $E(k)$ as we vary $\bar{\rho}$. For $\bar{\rho}\simeq 0.4$, the spectrum follows $E(k)\sim k^{-5/3}$, which steepens to $E(k)\sim k^{-5/2}$ as the density increases. This transition is observed for both types of swimmers. Although the exponent $-5/3$ is reminiscent of the exponent for the inverse-cascade scaling regime in high-Reynolds-number 2D fluid turbulence [5, 4] and in low-Reynolds-number active-scalar turbulence [26], there are significant differences between our spectra and their 2D-fluid-turbulence counterparts [see, e.g., Ref. [5, 4]]. In brief, a conventional 2D fluid yields statistically homogeneous and isotropic turbulence, if it is forced sufficiently strongly at a length scale $l_{f}$; the resulting energy spectrum displays two spectral regimes – the first with an inverse cascade of energy and the second with a forward cascade of enstrophy (i.e., the mean-square vorticity); in the inverse-cascade regime $E(k)\sim k^{-5/3}$, whereas, in the forward-cascade regime, $E(k)\sim k^{-\varpi}$, with $\varpi=3$ in the absence of friction [3, 4]; the crossover between these two regimes occurs at a wavenumber $k_{f}\simeq 2\pi/l_{f}$. Energy spectra for bacterial turbulence are also markedly different from those that we find for algal turbulence; e.g., the simple Toner-Tu-Swift-Hohenberg (TTSH) model, which has been used to model bacterial turbulence in Bacillus subtilis and Escherichia coli, yields a low-$k$ power-law regime in $E(k)$, but with an activity-dependent spectral exponent [10] that saturates eventually to a value $\simeq-3/2$ at large activity [27].
The density spectrum $\Phi(k)\sim k^{1/2}$ at low wave numbers for low densities; and it approaches a constant for large densities, as shown in Fig. 2.
@@
Figures 3(a) and (b) [blue panel] present PDFs (denoted generically by $\mathcal{P}$) of (a) $\bm{u}_{x}$ and (b) the Okubo-Weiss parameter [28, 29, 5] $\Lambda$ [see Materials and Methods, Eq. (10)] of WT cell suspensions. Their counterparts for the mbo2 mutant are given, respectively, in Figs. 3(c) and (d) [beige panel]. The PDFs for the wild-type system are qualitatively similar to their counterparts in the mbo2-mutant systems; however, quantitative differences do appear, especially in the PDFs for $\Lambda$. Moreover, $\mathcal{P}(\Lambda)$ shows more dependence on $\bar{\rho}$ for the WT cells than it does for the mbo2 variant.
Note that the PDFs of $\bm{u}_{x}$ are markedly different from those that are obtained for conventional fluid turbulence [see, e.g., Ref. [5]] and bacterial turbulence in dense suspensions of Bacillus subtilis [see, e.g., Ref. [9]], both of which are Gaussian. By contrast, the PDFs of $\bm{u}_{x}$, in Figs. 3(a) and (c), are distinctly non-Gaussian (we give fits to sums of exponentials and compressed exponentials in the Supplementary Information in Fig. 2). Furthermore, $\mathcal{P}(\Lambda)$ in the algal-turbulence state of our dense suspensions of C. reinhardtii cells is significantly different from its counterparts in 2D fluid turbulence  [30] and bacterial turbulence  [13].
Figure 4(a) [blue panel] displays PDFs ($\mathcal{P}$) of the longitudinal velocity increments $\delta\bm{u}_{\parallel}$(l) of WT cell suspensions  [5] [see Materials and Methods, Eq. (9)] for different values of $l$; Fig. 4(c) [beige panel] gives these PDFs for the mbo2 mutant. The $l$-dependence of these velocity-increment PDFs is a fingerprint of intermittency, which we quantify by the flatness $F_{4}$ that is related to the fourth- and second-order velocity structure functions [see Materials and Methods, Eq. (11)]. We plot the flatness, $F_{4}$ for both WT (versus $l/a_{w}$) and mbo2 (versus $l/a_{m}$) variants, in Figs. 4 (b) and 4 (d), respectively. Both these plots show distinct deviations from the Gaussian value $F_{4}^{G}=3$; this deviation increases as $l/a_{w}$ (or $l/a_{m}$) decreases, a clear manifestation of small-scale intermittency in the algal turbulence we consider  [^1] Signatures of small-scale intermittency are often uncovered by such plots of flatness or the hyperflatness $F_{6}\equiv S_{6}/(S_{2})^{3}$ in fluid and other forms of turbulence; see, e.g., Refs. [1, 46, 47, 5].. Furthermore, $F_{4}$ shows more dependence on $\bar{\rho}$ for the mbo2 mutant than it does for the WT variant, even though the error bars are large.

### v3

# First observation of turbulence in dense algal suspensions

@@ Abstract
Active turbulence arises typically in systems ranging from microorganisms and biopolymers to synthetic colloids, where chaotic flows are closely associated with motile topological defects in collectively swarming suspensions. Here, we report the first experimental observation of turbulence in a fundamentally different class of systems: dense monolayers of motile unicellular alga Chlamydomonas reinhardtii that exhibit neither orientational order nor topological defects. Nevertheless, the system displays rich spatiotemporal flow patterns with pronounced small-scale intermittency. We uncover strongly non-Gaussian velocity distribution, a feature distinct from both bacterial and classical fluid turbulence. Furthermore, we observe power-law regimes in the kinetic energy spectra, characterized by unique scaling exponents. Not only do our results provide compelling evidence for active turbulence in systems devoid of nematic or polar structures, but they also challenge current theoretical models. Our work opens new avenues for understanding emergent dynamics in active-matter systems and suggests intriguing biological implications, including enhanced mixing and transport in dense cell suspensions.
@@ I Uncovering Algal turbulence
Turbulence, which is ubiquitous from astrophysical to cellular scales Frisch-CUP; biskamp2003magnetohydrodynamic; pandit2009statistical; boffetta2012two; pandit2017overview; benzi2023lectures; alert2020universal, continues to provide new and exciting challenges for physicists, engineers, and mathematicians. Nonequilibrium turbulence-like states, which have been found over the last decade or so in dense bacterial suspensions wensink2012meso; dunkel2013fluid; bratanov2015new; wang2017bactmulitfrac; alert2021; KiranPRF2023; padhan2025suppression, epithelial-cell monolayers Lin2021energetics; blanch2018turbulent; Doostmohammadi2015-re, sperm suspensions creppy2015turbulence, and microtubule-motor mixtures Sanchez2012-or; guillamat2017taming; martinez2021scaling; maryshev2019dry provide important recent examples of new types of turbulence seen in biological systems. This has been christened active turbulence because such suspensions are active fluids in which energy is injected into the fluid not by an external force, as in conventional fluid turbulence, but by the conversion of chemical sources of energy to kinetic energy by the constituents in the suspension alert2021. So far, all the different active fluids, biological or synthetic, that show turbulence also have swarming motion of the constituents that require nematic or polar order parameters pandit2025particles; giomi2015geometry; Yashunsky2024-hr. Recent theoretical studies have shown that, surprisingly, even active-scalar fluids can also display turbulent states with rich phenomenology tiribocchi2015active; nardini2017entropy; pandit2025particles; padhan2024novel; padhan2023activity; Keta2024emerging.
Here, we present the first experimental study of emergent active turbulence and intermittency in non-swarming dense algal suspensions. In particular, we carry out experiments on suspensions of two types of Chlamydomonas reinhardtii (henceforth, C. reinhardtii), the wild type CC1690 (WT) and the mutant CC2377 (mbo2). We then characterise the statistical properties of this algal turbulence using measures that are employed to study conventional fluid turbulence.
We obtain kinetic-energy and algal-concentration spectra, longitudinal-velocity structure functions Frisch-CUP; pandit2009statistical; boffetta2012two; pandit2017overview; BuariaPRL2023, probability distribution functions (PDFs) of velocity components, length-scale dependent longitudinal-velocity increments, and the Okubo-Weiss parameter $\Lambda$, which distinguishes between vortical and extensional regions in a flow. These spectra display power-law scaling regions as a function of the wavenumber $k$; and they indicate that kinetic-energy and concentration fluctuations are spread over a wide range of spatial scales; these are clear signatures of turbulence. However, the power-law exponents that characterise these scaling regions are distinctly different from their fluid-turbulence and bacterial-turbulence counterparts Frisch-CUP; pandit2009statistical; boffetta2012two; pandit2017overview; alert2021, as are the PDFs and structure functions mentioned above. Thus, our investigations uncover a new type of active turbulence, which we call algal turbulence, and whose special statistical properties we elucidate below. Earlier experimental and theoretical studies leptos2009dynamics; thiffeault2015distribution; gollub2011enhancement have investigated enhanced tracer diffusion in dilute suspensions of swimming eukaryotic swimmers like C. reinhardtii. However, this enhanced tracer diffusion is more akin to Lagrangian chaos or passive-scalar intermittency in simple flows dombre1986chaotic; warhaft2000passive; falkovich2001particles than to the fully developed algal turbulence we have uncovered here.
As we increase the concentration of C. reinhardtii cells, the temporal evolution of our dense algal suspensions slows down and the intensity of algal turbulence in our systems decreases. Therefore, it behooves us to explore whether our algal systems cross over from algal turbulence to an active algal glass. To examine this possibility for the cells in our suspensions, we calculate several quantities that are used to characterize slow dynamics in active glasses sadhukhan2024perspective; Mandal2016-go; Henkes2011-hn, including (a) the mean-square displacement (MSD), (b) the self-intermediate scattering function $F_{s}(k,t)$, (c) the overlap function $Q(t)$, and (d) the four-point correlation function $\chi_{4}(t)$, which characterizes the dynamic heterogeneity in a glassy system. Our systems exhibit dynamic heterogeneity and distinct slowing down as $\bar{\rho}$ increases, but they do not display all the characteristic properties of an active glass.
@@
Figure 1: Time-lapse images of the motion of an isolated (a) wild-type (WT) and (b) mbo2 C. reinhardtii cell [scale bar, $10$ microns]. The flagella of the cells in these images have been highlighted manually for clarity. Experimentally measured beat-averaged flow-fields of isolated (c) WT and (d) mbo2 C. reinhardtii cells; the arrow indicates the direction of motion. (e) Interpolated velocity-vectors, overlaid on the image of a dense suspension of WT C. reinhardtii cells. The supplementary video V7 shows its spatiotemporal evolution.
II Statistical characterisation of dense algal suspensions
@@
The swimmers collectively churn the fluid, which leads to an emergent chaotic flow, with more complex vortical fields than those generated by individual swimmers. We characterise this complexity systematically using statistical measures that are employed to analyse statistically homogeneous and isotropic turbulence in fluids (see below). Figure 1(e) shows an optical image of the swimmers, at a representative time, overlaid with velocity vectors of the flow field, whose magnitude is given by the colour that goes from blue (low speed) to red (high speed). We note that the C. reinhardtii cells have a low aspect ratio that is very close to 1 and is much smaller than in bacteria like B. subtilis, whose aspect ratio varies from 4 to 8  ilkanaiv2017effect. Consequently, there is no nematic phase in the absence of activity, and the C. reinhardtii cells lack any intrinsic alignment because of their shape. This is also borne out by the rapid decay of the spatial velocity-velocity correlation function in our systems (Appendix D).
Figure 2: (a) Optical image of a suspension of WT C. reinhardtii cells, with average density $\bar{\rho}=0.63$ [scale bar, $50$ microns]; pseudocolor plots of (b) the density and (c) the vorticity $\omega$ fields, for the image shown in (a). Panels (d), (e), and (f) are the mbo2 C. reinhardtii counterparts of (a), (b), and (c), respectively.
@@
Figure 3: (a)-(b) Log-log plots of the energy spectra $E(k)$, compensated by different powers of $k$ and plotted versus $ka_{w}$ for WT cells with $\bar{\rho}=0.47,0.61,0.74$; (c) Log-log plots of the concentration spectra $\Phi(k)$, plotted versus $ka_{w}$. Panels (d), (e), and (f) are the mbo2 C. reinhardtii counterparts of (a), (b), and (c), respectively. Dark-gray shading indicates scaling regions. Our data are averaged over different samples with similar values of $\bar{\rho}$ (within $\simeq 10\%$ of the mean); the error-bars denote one-standard-deviation ($\varsigma$), i.e., $\pm\varsigma(E(k)/E_{0})(ka_{w})^{\frac{-1}{4}})$, $\pm\varsigma((E(k)/E_{0})(ka_{w})^{\frac{9}{2}})$ and $\pm\varsigma(\Phi(k)$. We define $E_{0}=\sum E(k)$.
We monitor the distributions of cells in dense suspensions of C. reinhardtii, and show these in Fig. 2 for WT cells (here and henceforth in a blue panel) and the mbo2 mutant (here and henceforth in a beige panel). Figure. 2(a) shows a microscopic snapshot of a WT suspension with an average density $\bar{\rho}=0.63$, at a representative time. At the same instant, we show a pseudocolor plot of the density field $\rho$, which lies between 0 (no cell present) and $1$ (complete coverage with the cell bodies) [Fig. 2(b)] and the vorticity-field $\omega$ computed from the velocity field $\bm{u}$ [Fig. 2(c)]. Figures 2 (d), (e), and (f) are, respectively, the counterparts of Figs. 2 (a), (b), and (c) for the mbo2 mutant. Supplementary videos V1-V4 are raw experimental videos that show the spatiotemporal evolution of WT and mbo2 cell suspensions. Supplementary videos V5 and V6 show the spatiotemporal evolution of the density and the vorticity fields (WT cells, $\bar{\rho}=0.50$). These plots and the videos show that dense suspensions of both WT and mbo2 C. reinhardtii cells show emergent nonequilibrium states that appear to display spatiotemporal chaos and a type of turbulence. We quantify the statistical properties of these states below.
In Fig. 3 we present log-log plots versus the wavenumber $k$ of the energy spectra $E(k)$ [Eqs. (2) and (3) in Subsection IV.5] and the concentration spectra $\Phi(k)$ [Eqs. (1),(3) in Subsection IV.5] for the wild-type (blue panels) and mbo2 mutant (beige panels). We use $a_{w}$ ($a_{m}$), the mean wild-type (mbo2-mutant) cell size (Appendix A) to non-dimensionalize $k$. We also compensate the energy spectra with powers of $k$ to uncover different power-law regimes. The energy spectra scales as $E(k)\sim k^{1/4}$ in the small-$k$ limit and $E(k)\sim k^{-9/2}$ in the large-$k$ limit for all $\bar{\rho}$, and for both types of swimmers. But the density spectrum $\Phi(k)$ of WT and mbo2 cells differ from each other. This is because the spatial distribution of WT cells is different from that of the mbo2 mutant (Appendix K and Appendix L); and this distribution of mbo2 cells (at $\bar{\rho}=0.50$), which are extensile swimmers, is somewhat reminiscent of phase separation  padhan2024novel.
@@
There are significant differences between our spectra and their 2D-fluid-turbulence counterparts pandit2017overview; boffetta2012two. In brief, a conventional 2D fluid yields statistically homogeneous and isotropic turbulence, if it is forced sufficiently strongly at a length scale $l_{f}$; the resulting energy spectrum displays two spectral regimes – the first with an inverse cascade of energy and the second with a forward cascade of enstrophy (i.e., the mean-square vorticity); in the inverse-cascade regime $E(k)\sim k^{-5/3}$, whereas, in the forward-cascade regime, $E(k)\sim k^{-\varpi}$, with $\varpi=3$ in the absence of friction pandit2009statistical; boffetta2012two; the crossover between these two regimes occurs at a wavenumber $k_{f}\simeq 2\pi/l_{f}$. Energy spectra for bacterial turbulence are also markedly different from those that we find for algal turbulence; e.g., the simple Toner-Tu-Swift-Hohenberg (TTSH) model, which has been used to model bacterial turbulence in Bacillus subtilis and Escherichia coli, yields a low-$k$ power-law regime in $E(k)$, but with an activity-dependent spectral exponent bratanov2015new that saturates eventually to a value $\simeq-3/2$ at large activity mukherjee2023intermittency.
We compute two natural characteristic length scales $L_{E}=2\pi\frac{\sum E(k)/k}{\sum E(k)}$ and $L_{\phi}=2\pi\frac{\sum\phi(k)}{\sum k\phi(k)}$ from the energy spectra and the density spectra (Appendix EPerlekar2017two; Dhar1997some. $L_{E}$ stays constant around $50\mu m$ and $L_{\phi}$ stays constant around $25\mu m$ for WT cells ($\bar{\rho}=0.50$). This suggests that the algal system is in a statistically steady state. Also, $L_{\phi}$ seems to increase monotonically with average density $\bar{\rho}$ for WT cells; but this behavior is not seen in mbo2 cells.
@@
Figure 4 [blue panel] presents PDFs ($\mathcal{P}$) of (a) $\bm{u}_{x}$ and (b) the Okubo-Weiss parameter $\Lambda$ of WT cell suspensions [Materials and Methods, Eq. (5weiss1991dynamics; okubo1970horizontal; pandit2017overview]. Their counterparts for the mbo2 mutant are given, respectively, in Figs. 4(c) and (d) [beige panel]. The skewness for WT cells of $\bar{\rho}\sim 0.47,0.61$, and $0.74$ are $0.29,0.40$, and 0.70-0.70, respectively. For the mbo2 cells with $\bar{\rho}\sim 0.63,0.69$, and $0.76$, the corresponding skewness values are 0.17,0.320.17,-0.32, and 0.70-0.70, respectively. The PDFs for the wild-type system are qualitatively similar to their counterparts in the mbo2-mutant systems.
Figure 4: Semilog plots of different probability distribution functions (PDFs): PDFs of (a) the $x$-component of the velocity field $\bm{u}$ and (b) the Okubo-Weiss parameter $\Lambda$ [Eq. (5)] for WT C. reinhardtii cells and different mean densities $\bar{\rho}$. Panels (c) and (d) are the mbo2 C. reinhardtii counterparts of (a) and (b), respectively.
Figure 5: Semilog plots, for WT C. reinhardtii cells, of (a) the PDFs of longitudinal-velocity increments of the velocity field $\bm{u}$, for different mean densities $\bar{\rho}$ and the separation $l/a_{w}=5.37$. (b) the flatness $F_{4}$ [Eq. (6)] versus $l/a_{w}$ for different values of $\bar{\rho}$. The plots in (c) and (d) are the mbo2-mutant counterparts of those in (a) and (b), respectively. Our data are averaged over different samples with similar values of $\bar{\rho}$ (within $\simeq 10\%$ of the mean); the error bars denote $\pm\varsigma(F_{4})$.